MEASUREMENTS OF GROUND STATE OXYGEN, WATER VAPOR AND TEMPERATURE IN CHEMICAL OXYGEN-IODINE LASER BY TI:SAPPHIRE LASER-BASED INTRACAVITY LASER SPECTROSCOPY

\underline{M. P. FROLOV}, Yu. P. PODMARKOV ^{a}, Lebedev Physical Institute, Leninsky pr. 53, Moscow, 117924 Russia.

The ground state oxygen and water vapor are important species in chemical oxygen-iodine laser. The detectivity required in the chemical oxygen-iodine laser device is about 0.03 Torr for both H_{2}O and O_{2}(X). We have shown that these parameters and the temperature of chemical oxygen-iodine laser medium can be measured by intracavity laser spectroscopy based on simple standing wave Ti:sapphire laser. Using 120-μs laser generation time, we have demonstrated the detection sensitivities to be 6 · 10^{1}4 and 3 · 10^{1}4 cm^{−3} for H_{2}O and O_{2}(X), respectively. The oxygen concentration measurement error does not exceed 5% in the range 2 · 10^{1}5–10^{1}6 cm^{−3}. The temperature measurement error does not exceed 25K in the range 200-500K.

aSupport by Russian Foundation for Basic Research grant is gratefully acknowledged